# Measuring Physical-World Privacy Awareness of Large Language Models: An Evaluation Benchmark

### v2

@@ Appendix D Standard Deviation of Results
Appendix E Full Results
@@
E.1 Complete Tier 1 Results
@@
E.2 Complete Tier 2 Results
@@
E.3 Complete Results Table
@@
Appendix F Case Study Details for Tier 1
@@
Appendix G Case Study Details for Tier 2
@@
Appendix H Case Study Details for Tier 3
@@
Appendix I Case Study Details for Tier 4
@@
Appendix J Tier 1 Spatial Evaluation Metrics
@@
Appendix K Example Input of Each Tier
@@
K.1 Example PDDL File for Tier 1 Scenario
PDDL file representations are widely used in robotics and embodied AI (Li et al., 2025; Shridhar et al., 2020) to represent spatial relationships and object states. Below,  K.1 , are example PDDL files for a Tier 1 scenario from our evaluation set. It provides structural representations for the spatial configuration of objects and their states in the environment, beyond simple natural language conversations, descriptions or QA.
@@
K.2 Example Input for Tier 2 Scenario
In this tier, we use widely adopted structural spatial representations PDDL files in embodied agent frameworks (Li et al., 2025; Shridhar et al., 2020) and sensory-like observations to mimic the input an embodied agent would receive. In practice, multimodal modules would convert raw sensory data (e.g., images, point clouds) into parsed summaries or communication (Aissi et al., 2025; Park et al., 2023), we use action return lists to simulate this process. Below,  K.2 and  K.2, is an example input for a Tier 2 scenario from our evaluation set, such input goes beyond simple natural language conversations to simulate embodied agent input.
@@
K.3 Example Input for Tier 3 Scenario
In this tier, we also use widely adopted structural spatial representations PDDL files in embodied agent frameworks (Li et al., 2025; Shridhar et al., 2020) and sensory-like observations to mimic the input an embodied agent would receive. In practice, multimodal modules would convert raw sensory data (e.g., images, point clouds) into parsed summaries or communication (Aissi et al., 2025; Park et al., 2023), we use action return lists to simulate this process. Below,  K.3 and  K.3, is an example input for a Tier 3 scenario from our evaluation set, such input goes beyond simple natural language conversations to simulate embodied agent input.
@@
K.4 Example Input for Tier 4 Scenario
In this tier, we also use widely adopted structural spatial representations PDDL files in embodied agent frameworks (Li et al., 2025; Shridhar et al., 2020) and sensory-like observations to mimic the input an embodied agent would receive. In practice, multimodal modules would convert raw sensory data (e.g., images, point clouds) into parsed summaries or communication (Aissi et al., 2025; Park et al., 2023), we use action return lists to simulate this process. Below,  K.4 and  K.4, is an example input for a Tier 4 scenario from our evaluation set, such input goes beyond simple natural language conversations to simulate embodied agent input.
@@
Appendix L Action List PDDL File
@@
L.1 Part of the Action List PDDL File

### v3

@@ Appendix F Standard Deviation of Results
Appendix G Full Results
@@
G.1 Complete Tier 1 Results
@@
G.2 Complete Tier 2 Results
@@
G.3 Complete Results Table
@@
Appendix H Case Study Details for Tier 1
@@
Appendix I Case Study Details for Tier 2
@@
Appendix J Case Study Details for Tier 3
@@
Appendix K Case Study Details for Tier 4
@@
Appendix L Tier 1 Spatial Evaluation Metrics
@@
Appendix M Example Input of Each Tier
@@
M.1 Example PDDL File for Tier 1 Scenario
PDDL file representations are widely used in robotics and embodied AI (Li et al., 2025; Shridhar et al., 2020) to represent spatial relationships and object states. Below,  M.1 , are example PDDL files for a Tier 1 scenario from our evaluation set. It provides structural representations for the spatial configuration of objects and their states in the environment, beyond simple natural language conversations, descriptions or QA.
@@
M.2 Example Input for Tier 2 Scenario
In this tier, we use widely adopted structural spatial representations PDDL files in embodied agent frameworks (Li et al., 2025; Shridhar et al., 2020) and sensory-like observations to mimic the input an embodied agent would receive. In practice, multimodal modules would convert raw sensory data (e.g., images, point clouds) into parsed summaries or communication (Aissi et al., 2025; Park et al., 2023), we use action return lists to simulate this process. Below,  M.2 and  M.2, is an example input for a Tier 2 scenario from our evaluation set, such input goes beyond simple natural language conversations to simulate embodied agent input.
@@
M.3 Example Input for Tier 3 Scenario
In this tier, we also use widely adopted structural spatial representations PDDL files in embodied agent frameworks (Li et al., 2025; Shridhar et al., 2020) and sensory-like observations to mimic the input an embodied agent would receive. In practice, multimodal modules would convert raw sensory data (e.g., images, point clouds) into parsed summaries or communication (Aissi et al., 2025; Park et al., 2023), we use action return lists to simulate this process. Below,  M.3 and  M.3, is an example input for a Tier 3 scenario from our evaluation set, such input goes beyond simple natural language conversations to simulate embodied agent input.
@@
M.4 Example Input for Tier 4 Scenario
In this tier, we also use widely adopted structural spatial representations PDDL files in embodied agent frameworks (Li et al., 2025; Shridhar et al., 2020) and sensory-like observations to mimic the input an embodied agent would receive. In practice, multimodal modules would convert raw sensory data (e.g., images, point clouds) into parsed summaries or communication (Aissi et al., 2025; Park et al., 2023), we use action return lists to simulate this process. Below,  M.4 and  M.4, is an example input for a Tier 4 scenario from our evaluation set, such input goes beyond simple natural language conversations to simulate embodied agent input.
@@
Appendix N Action List PDDL File
@@
N.1 Part of the Action List PDDL File